OpenStudy (anonymous):

Which ordered pair is a solution of the system of equations y = x + 3 and y = x2 − x?

OpenStudy (mertsj):

Just plug them in and see which makes a true sentence.

OpenStudy (anonymous):

these are the choices (1) (6,9) (3) (3,−1) (2) (3,6) (4) (2,5)

OpenStudy (anonymous):

let one variable = 0 and solve for the other i believe there is multiple answers for this question however if you have a list of pairs put them in

OpenStudy (anonymous):

alright so for each pair (x,y) ... put it into the equation if it's true for both equations, it is a solution pair to the system of equations

OpenStudy (mertsj):

That is correct.

OpenStudy (pfenn1):

Or, since both equations are equal to y, you can set them equal to one another like so\[x^2-x=x+3\]Subtract x + 3 from both sides to get\[x^2-2x-3=0\]Either factor or use the quadratic formula to solve.